Distribution of Cash Upon Liquidation Hewitt and Patel are partners, sharing gains and losses equally. They decide to terminate their partnership. Prior to realization, their capital balances are $42,000 and $28,000, respectively. After all noncash assets are sold and all liabilities are paid, there is a cash balance of $53,000. a. What is the amount of a gain or loss on realization?
Answer #1

Since no information has been given about the cash balance appearing in the balance sheet before the liquidation, it can be assumed that there was no cash appearing in the balance sheet.

The cash balance after selling the non cash assets and liabilities is $53,000

The capital accounts of the partners before the realization represent the difference between the assets and liabilities.

The total of capital balances before the realization was 42000 + 28000 = $70000.

Loss on realization = Capital account balance - Net cash available

Hence, there was a loss in realization = 70000 - 53000 = $17000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.